OpenStudy (anonymous):

@undeadknight26 When using the Babylonian method to approximate square root of 5 if you start with a guess of 2, then what will be your next guess? A. 2 B. 2.5 C. 2.25 D. 3

OpenStudy (anonymous):

can you check my other questions i have awnswered them all exept this one

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

The Babylonian method works as follows: Given an approximation xn to the square-root of N, the next approximation is obtained as xn+1 = (xn + N/xn)/2 For example, if the initial approximation of sqrt (10) is 3, then the next approximation is (3+10/3)/2 = 19/6

OpenStudy (anonymous):

Arrange the following fractions in order from least to greatest. 1/3 2/5 3/8 i got 1/3 3/8 2/5

OpenStudy (anonymous):

Find a number that is between 2/7and 4/11 i got 24/77

OpenStudy (anonymous):

Simplify completely. sqrt20y^8 i got 2y^4sqrt 5

OpenStudy (anonymous):

Between which two consecutive integers does sqrt89 lie? i got 9 and 10

OpenStudy (anonymous):

Simplify sqrt 0.25 i got 0.5

OpenStudy (anonymous):

@mathmate @undeadknight26

undeadknight26 (undeadknight26):

No idea sorry mate...

OpenStudy (anonymous):

what about the other 5

OpenStudy (anonymous):

so mathmate is the first one b?

OpenStudy (anonymous):

nvm i got 5/6 right so i got 83% but tks